# Versatile Motion-Language Models for Multi-turn Interactive Agents

### v2

@@ 5.4 Traditional Motion Related Tasks
The results in Table 7 support our hypothesis that utilizing the Inter-$\text{MT}^{2}$ dataset enhances the model’s performance in traditional motion tasks like motion-to-text (M2T), text-to-motion (T2M), and reaction generation. The first row (“Real”) shows retrieval accuracy, and FID scores from the dataset labels. Note that both VIM w/o Inter-$\text{MT}^{2}$ and MotionGPT were trained on all of these tasks for fair comparison. Comparing the VIM w/o Inter-$\text{MT}^{2}$ to the version trained with Inter-$\text{MT}^{2}$ (“Ours”), we see improvements across all tasks. In M2T, Top-3 retrieval accuracy rose from 0.894 to 0.901. For T2M, Top-3 retrieval accuracy increased from 0.561 to 0.568, with FID dropping from 0.082 to 0.059, indicating better motion generation. In reaction generation, MPJPE dropped from 0.984 to 0.691, and FID from 0.031 to 0.019, confirming that multi-turn datasets improve motion comprehension and generation. Using the Inter-$\text{MT}^{2}$ dataset provides diverse, context-rich examples, helping the model learn more nuanced relationships between text and motion. Additionally, incorporating Inter-$\text{MT}^{2}$ in $\text{MotionGPT}^{*}$, denoted as $\text{MotionGPT}^{*}_{I}$, improved retrieval precision accuracy for M2T and T2M tasks, and joint position error in reaction generation.
5.5 Ablation Studies on Motion Tokenizer
Table 4: Ablation Studies on motion tokenizer.
Methods Reasoning Editing M2T T2M Reaction Gen.
VIM-VQ 5.004 4.256 6.915 0.892 0.128 0.861 0.601 0.101 1.109 0.055
VIM (Ours) 5.252 4.511 6.981 0.758 0.064 0.901 0.568 0.059 0.691 0.019
We conducted ablation studies comparing the VQ-VAE-based model with our RQ-VAE-based approach, as shown in Table 4. The RQ-VAE-based motion tokenizer outperformed the VQ-VAE model in motion reasoning tasks, achieving higher scores in coherence, alignment, and naturalness. This improvement is attributed to reduced information loss, allowing our model to capture finer motion details while also enhancing its motion-to-text retrieval precision. For generation and editing tasks, the VQ-VAE model achieved slightly better text-to-motion retrieval accuracy but performed worse in FID and MPJPE across editing, reaction generation, and T2M tasks, indicating degraded motion quality and less precise motion details. In contrast, our approach reduced MPJPE by 0.055 for reaction generation, preserving joint dynamics and producing more realistic and natural motions. VQ-VAE’s limitations are especially problematic for modeling interactive motions, where precise relative positioning is crucial, making its information loss and reconstruction quality more evident.
6 Conclusion and Discussions
Conclusion
@@
In this paper, we introduced VIM, a versatile motion-language model designed to model, understand, and reason about interactive motions. We outlined its architecture and provided detailed training strategies to create a unified framework integrating large language models with interactive motion modality. To enhance the model’s reasoning capabilities and versatility, we presented a specialized dataset, Inter-$\text{MT}^{2}$, which incorporates a variety of reasoning tasks set within multi-turn conversations centered on interactive motions. Our experiments demonstrated VIM’s ability to effectively follow instructions, edit motions, and reason about interactive motions.
Limitations and Impact Statement
There are several limitations that warrant attention. First, the model’s expressiveness remains limited when handling complex or previously unseen actions, indicating a need for further diverse motion source data in its ability to generalize across diverse motion scenarios. Second, the sequence length becomes excessively long as we flatten the residual motion tokens, which can impact efficiency and computational resources. Leveraging additional transformer models to predict the residual token can reduce this work. Lastly, our method faces challenges in personalization and interpretability, as motion is inherently ambiguous and users may interpret the same motion in different ways. Addressing this issue will require incorporating more tailored approaches that adapt to individual user preferences and expectations through further human-in-the-loop feedback and refinement processes. In terms of broader impact, VIM opens up new possibilities for interactive motion modeling and understanding in AI, potentially benefiting fields like robotics, virtual environments, and human-computer interaction. However, careful consideration of ethical concerns, such as misinterpretation of motions or unintended behavioral biases, is crucial as the model evolves.
Appendix A Appendix
@@
This appendix provides a comprehensive set of supplementary materials that reinforce the main findings of the research. It covers key areas such as motion representation (Sec. A.1) and Inter-$\text{MT}^{2}$ dataset sample visualization (Sec. A.2), with accompanying dataset statistics (Sec. A.3). In-depth task explanations are included (Sec. A.4), alongside ablation studies that examine various pretraining methods (Sec. A.5). The appendix also contains qualitative results (Sec. A.6) and thorough explanations of two-stage baselines (Sec. A.8). Additionally, it provides template forms for pre-training and instruction tuning (Sec. A.9). We also report implementation details for MotionGPT (Sec. A.7), with implementation details for the proposed method (Sec. A.10), detailed metrics explanation (Sec. A.11), protocols for user subject studies (Sec. A.12) focused on motion editing, prompts for data collection within the dataset (Sec. A.13), and guidelines for LLM-assisted evaluation processes (Sec. A.14).
A.1 Motion Representation and Motion Token Representation
@@
For two persons $a$ and $b$, we denote the interactive motion as $\{\mathbf{m}_{a},\mathbf{m}_{b}\}$, following non-canonical representation from Liang et al. (2024). Each timestep of the motion $\mathbf{m}^{i}=[\mathbf{j}_{g}^{p},\mathbf{j}_{g}^{v},\mathbf{j}^{r},\mathbf{c}^{f}]$ is composed of global joint positions $\mathbf{j}_{g}^{p}\in\mathbb{R}^{3N_{j}}$, global joint velocities $\mathbf{j}_{g}^{v}\in\mathbb{R}^{3N_{j}}$, 6D representation of local rotations $\mathbf{j}^{r}\in\mathbb{R}^{6N_{j}}$, with the number of joints $N_{j}$, and binary ground contact features $\mathbf{c}^{f}\in\mathbb{R}^{4}$. This non-canonical representation is applied for both interactive motions and single-person motions. All the motions are represented in an SMPL-X (Pavlakos et al., 2019) format.
@@
where <motion_token_start>, <motion_token_a_start>, <motion_token_b_start>, <motion_token_a_end>, <motion_token_b_end>, and <motion_token_end> is a special token added to the unified vocabulary.
A.2 Data Sample Visualization
The samples from the synthesized dataset, Inter-$\text{MT}^{2}$, are illustrated in Figure 7.
Figure 7: Sample from Inter-$\text{MT}^{2}$ datset. The left column visualizes samples of motion editing, and the right column shows the examples from motion reasoning task.
A.3 Inter-$\text{MT}^{2}$ Statistics
Table 5: Statistics on Inter-$\text{MT}^{2}$.
Total Train Val. Test
# of Samples 82736 66194 4141 12401
# of Motions 317749 132388 8282 24802
From Dataset 56395 50258 3142 2995
Synthesized 96676 82130 5140 9406
Table 6: Comparison of retrieval precision, motion diversity (Div.), and motion quality metrics (MMDist. and FID) across synthesized and source motions. The synthesized motion dataset (96K pairs) shows a Top 3 retrieval precision of 0.668, comparable to the InterGEN model’s precision (0.645) on the InterX+H dataset, indicating competitive text-to-motion matching quality.
Source # of pairs Retrieval Precision MMDist. Div. FID
Top1 Top2 Top3
InterX+H 18K 0.645 0.804 0.870 1.072 0.997 -
Synthesized Motion 96K 0.480 0.595 0.668 1.102 0.824 -
Model Dataset
InterGEN Liang et al. (2024) InterX+H 0.403 0.552 0.645 1.115 0.953 0.078
We collected 82K samples of multi-turn conversational data, each involving interactive motions. Of these, 30K samples focus on motion editing, 30K on reasoning about past or future scenarios, and 12K on story generation. Each sample includes four to eight conversation turns and two distinct motions. The dataset contains 96K motions generated using a text-to-motion diffusion model, while 56K motions come from the original source dataset. The train-validation-test set is randomly splitted by the ratio 0.8:0.05:0.15.
The quality of these motions is detailed in Table 6. From the generated caption from a large language model, we evaluate the text-motion matching score based on retrieval precision based on the feature space of retrieval models from Petrovich et al. (2023). This evaluates the accuracy of matching between texts and motions using Top 3 retrieval accuracy with a fixed batch of 32. The table’s first row shows the retrieval models’ performance, with a Top 3 retrieval precision of $0.870$. We found that the synthesized motions achieve a Top 3 retrieval precision of $0.668$, closely aligning with the reported precision of $0.645$ from the text-to-motion diffusion model (Liang et al. (2024)). This demonstrates that the synthesized motions maintain a high level of quality, making the dataset valuable and suitable for further training and development.
A.4 Detailed Task Explanations
Motion Editing
Motion Reasoning
A.5 Ablation Studies on Pretraining Method
Table 7: Ablation studies in pertaining stage for three motion-related tasks on InterX and Interhuman dataset.
@@ A.5 Ablation Studies on Pretraining Method
We conducted ablation studies on the pertaining method. All the baselines are pre-trained models, not including the fine-tuning stage. To evaluate the effectiveness of our pretraining approach, we conducted ablation studies comparing different methods on three motion-related tasks: Motion-to-Text (M2T), Text-to-Motion (T2M), and Reaction Generation. As shown in Table 7, we compared our proposed method, VIM, against MotionGPT and VIM-VQ, using the InterX (Xu et al., 2024a) and Interhuman (H) datasets (Liang et al., 2024). MotionGPT serves as a baseline with 248M trainable parameters, achieving a retrieval Top3 score of 0.518 in M2T and 0.280 in T2M, with corresponding FID scores of 0.178 and 1.338 for T2M and Reaction Generation, respectively. VIM-VQ, with 726M parameters, improves the M2T retrieval Top3 to 0.709 and T2M retrieval Top3 to 0.511, while maintaining competitive FID scores.
@@
Our method, VIM, further enhances performance by achieving a retrieval Top3 of 0.721 in M2T and reducing the T2M FID to 0.161, alongside an MPJPE of 1.494 and FID of 0.157 in Reaction Generation. Notably, when incorporating the additional MotionX (Lin et al., 2024) dataset, VIM achieves the highest M2T R Top3 of 0.729 and the lowest FID scores of 0.172 in T2M and 0.131 in Reaction Generation, demonstrating the substantial benefits of our comprehensive pretraining strategy. These results indicate that our approach not only outperforms existing models in generating accurate and high-quality motions but also effectively leverages additional data to enhance interactive motion understanding and generation. The ablation studies highlight the critical role of our pretraining methodology and the integration of diverse datasets in achieving superior performance across multiple interactive tasks.
A.6 Qualitative Results
We visualize our result gallery on motion editing in Figure 8 and on motion reasoning in Figure 9. Furthermore, the results for motion-to-text (Figure 10), text-to-motion (Figure 11), and reaction generation (Figure 12) are demonstrated.
Figure 8: Generated motion from source and editing command.
Figure 9: Generated responses based on the previous conversations for motion reasoning task.
Figure 10: Motion-to-text results. The blue part is generated motion captions from source motions.
Figure 11: Text-to-motion results. The blue part is generated motions from the motion caption.
Figure 12: Reaction Generation. The input motion is orange, while the generated reactive motion is colored blue.
A.7 Implementation details for MotionGPT
For training MotionGPT (Jiang et al., 2023) in the interactive motion dataset, we have utilized the Flan-T5-base model (Chung et al., 2024) as a base large language model. We trained the model with Interhuman (Liang et al., 2024) and InterX (Xu et al., 2024a) dataset, with the non-canonical representation, same as the proposed method. Although scaling up the model can improve the performance, we conducted the experiment with the same base model as the original paper from MotionGPT (Jiang et al., 2023) and Motionchain (Jiang et al., 2024). The original paper reported that increasing the model size did not significantly improved the model’s performance.
A.8 Detailed Explanation about Two-stage Baselines
In Section 5.2 and Section 5.3, we have compared the proposed method with two-stage models. In particular, we have utilized TM2T (Guo et al., 2022) for the motion captioner and InterGEN (Liang et al., 2024) for the text-to-motion generation model.
Motion Editing
In the motion editing task, the two-stage approach first uses the motion-to-text (TM2T; Guo et al. (2022)) model to generate a caption from the source motion and append the editing command. Then, the text-to-motion (InterGen; Liang et al. (2024)) model produces the edited motion based on this caption and command. In particular, the input for text-to-motion model is ”[motion caption]. [editing command]”.
We first trained TM2T model with the InterHuman dataset (Liang et al., 2024)and the InterX Xu et al. (2024a) dataset, which we denote as ${}^{*}*$. The performance is shown in Table 8. The TM2T model shows substantial improvements over the baseline MotionGPT models across all evaluation metrics. Specifically, TM2T achieves Retriveal Precision scores of 0.413 (Top1), 0.589 (Top2), and 0.696 (Top3), along with BLEU, METEOR, and Rouge-L scores of 0.192, 0.386, and 0.395, respectively. These results indicate that the task-specific TM2T model effectively generates accurate and relevant motion captions, making it a reliable choice for motion editing tasks. Although there remains a performance gap compared to the proposed method, the TM2T model provides a robust foundation for generating moderate-quality motion captions.
Table 8: Motion-to-Text performance for TM2T
@@ A.8 Detailed Explanation about Two-stage Baselines
task-specific approach
TM2T 0.413 0.589 0.696 0.192 0.386 0.395
Table 9: Text-to-Motion performance for InterGEN

### v5

@@ 5.4 Traditional Motion Related Tasks
task-specific approach
[…]
$\text{MoMask}^{*}$ 0.612 0.066 1.602 0.112
ComMDM - 0.251 0.304 - -
InterGEN - 0.645 0.078 - -
unified approach
@@
In this section, we conduct comparison experiments on existing motion-relevant tasks, such as motion-to-text (M2T), text-to-motion (T2M), and reaction generation. The detailed results are in Table 4. The first row (“Real”) shows retrieval accuracy, and FID scores from the dataset labels. Note that both VIM w/o Inter-$\text{MT}^{2}$ and MotionGPT were trained on all of these tasks for fair comparison. The results confirm that incorporating Inter-$\text{MT}^{2}$ dataset enhances the model’s performance in traditional motion tasks, by comparing with VIM w/o Inter-$\text{MT}^{2}$.
For M2T, Top-3 retrieval accuracy improved from 0.894 (VIM w/o Inter-$\text{MT}^{2}$) to 0.901 (VIM). For T2M, it rose from 0.561 to 0.568, with FID dropping from 0.082 to 0.059, indicating better motion generation. For reaction generation, MPJPE decreased from 0.984 to 0.691 and FID from 0.031 to 0.019, highlighting the benefits of multi-turn datasets for motion comprehension and generation. We believe that Inter-$\text{MT}^{2}$ dataset provides diverse, context-rich examples, helping the model learn more nuanced relationships between text and motion.
In addition, we compared VIM against task-specific methods, each optimized individually per each task. Note that the methods marked with an asterisk (*) were originally designed for single-motion tasks and were trained on interactive motion data for our evaluation. VIM outperforms these specialized models in motion-to-text (M2T) and reaction generation tasks, achieving higher retrieval precision accuracy and lower MPJPE and FID scores. In the text-to-motion (T2M) task, VIM achieves comparable performance against the state-of-the-art task-specific models, including InterGEN and MoMask, highlighting the capability of VIM in generating high-quality interactive motions.
5.5 Discussion: Generating Multi-Human Motions
Figure 8: Expanding VIM to generate multiple human motions. For clarity, we simplified the incremental process, where VIM first generates two-person motion and adds a third-person to them.
Interestingly, the versatility of VIM allows it to generalize beyond two-person interactions without explicit fine-tuning on multi-person interactive data. Specifically, users can incrementally generate motions: first generating a two-person motion, then adding a third person’s motion conditioned on the existing context. By leveraging prior turns, VIM seamlessly integrates the new figure’s movements, as shown in Figure 8. Since our method is agnostic to the number of people, it can readily extend to groups or crowds, provided interactive multi-person data is available.
6 Conclusion
@@
In this paper, we introduced VIM, a versatile motion-language model designed to understand, generate, and reason about interactive motions. We presented the detailed architecture and training strategy of our unified framework, which integrates large language models with interactive motion modality. To further enhance the model’s reasoning capabilities and applicability, we presented a specialized dataset, Inter-$\text{MT}^{2}$, which incorporates a variety of reasoning tasks set within multi-turn conversations centered on interactive motions. Our comprehensive experiments demonstrated that VIM successfully handles instruction-following, motion editing, and motion reasoning tasks, highlighting its capability to effectively interpret and generate contextually accurate interactive motions.
Appendix
@@
This appendix provides a comprehensive set of supplementary materials that reinforce the main findings of the research. The appendix begins with motion representation and motion token representation (Sec.A), followed by ablation studies on the pretraining method (Sec.B), along with ablation studies on the motion tokenizer (Sec.C), the demonstration for expansion to multiple human ($\geq 3$) motion generation (Sec.D), and illustrations of the data collection pipeline (Sec. E). More detailed results for traditional motion-related tasks are presented (Sec.F), limitations (Sec.G), and implementation details for the proposed methods (Sec.H) and baselines models trained for interactive motions (Sec.I). Task explanations cover motion editing and reasoning (Sec.J), with implementation details of two-stage baselines (Sec.K). The evaluation metrics for traditional motion-related tasks are presented in Sec. L. Further sections include templates for pre-training and instruction tuning (Sec.M), data visualization and statistics (Sec.N, Sec.O), qualitative results (Sec.P), user study protocols (Sec.Q), and prompts for data collection and LLM-assisted evaluation (Sec.R, Sec. S).
A Motion Representation and Motion Token Representation
@@
For two persons $a$ and $b$, we denote the interactive motion as $\{\mathbf{m}_{a},\mathbf{m}_{b}\}$, following non-canonical representation from [22]. Each timestep of the motion $\mathbf{m}^{i}=[\mathbf{j}_{g}^{p},\mathbf{j}_{g}^{v},\mathbf{j}^{r},\mathbf{c}^{f}]$ is composed of global joint positions $\mathbf{j}_{g}^{p}\in\mathbb{R}^{3N_{j}}$, global joint velocities $\mathbf{j}_{g}^{v}\in\mathbb{R}^{3N_{j}}$, 6D representation of local rotations $\mathbf{j}^{r}\in\mathbb{R}^{6N_{j}}$, with the number of joints $N_{j}$, and binary ground contact features $\mathbf{c}^{f}\in\mathbb{R}^{4}$. This non-canonical representation is applied for both interactive motions and single-person motions. All the motions are represented in an SMPL-X [28] format.
@@
where <motion_token_start>, <motion_token_a_start>, <motion_token_b_start>, <motion_token_a_end>, <motion_token_b_end>, and <motion_token_end> is a special token added to the unified vocabulary. For modeling single-motion in pre-training we omitted the input string about motion_token_b.
B Ablation Studies on Pretraining Method
Table B.1: Ablation studies in pertaining stage for three motion-related tasks on InterX and Interhuman dataset.
@@ B Ablation Studies on Pretraining Method
We conducted ablation studies on the pertaining method. All the baselines are pre-trained models, not including the fine-tuning stage. To evaluate the effectiveness of our pretraining approach, we conducted ablation studies comparing different methods on three motion-related tasks: Motion-to-Text (M2T), Text-to-Motion (T2M), and Reaction Generation. As shown in Table B.1, we compared our proposed method, VIM, against MotionGPT and VIM-VQ, using the InterX [39] and Interhuman (H) datasets [22]. MotionGPT serves as a baseline with 248M trainable parameters, achieving a retrieval Top3 score of 0.518 in M2T and 0.280 in T2M, with corresponding FID scores of 0.178 and 1.338 for T2M and Reaction Generation, respectively. VIM-VQ, with 726M parameters, improves the M2T retrieval Top3 to 0.709 and T2M retrieval Top3 to 0.511, while maintaining competitive FID scores.
@@
Our method, VIM, further enhances performance by achieving a retrieval Top3 of 0.721 in M2T and reducing the T2M FID to 0.161, alongside an MPJPE of 1.494 and FID of 0.157 in Reaction Generation. Notably, when incorporating the additional MotionX [23] dataset, VIM achieves the highest M2T R Top3 of 0.729 and the lowest FID scores of 0.172 in T2M and 0.131 in Reaction Generation, demonstrating the substantial benefits of our comprehensive pretraining strategy. These results indicate that our approach not only outperforms existing models in generating accurate and high-quality motions but also effectively leverages additional data to enhance interactive motion understanding and generation. The ablation studies highlight the critical role of our pretraining methodology and the integration of diverse datasets in achieving superior performance across multiple interactive tasks.
C Ablation Studies on Motion Tokenizer
Table C.2: Ablation Studies on motion tokenizer base model. We compared VQ-VAE-based tokenizer and the RQ-VAE-based model.
Methods Reasoning Editing M2T T2M Reaction Gen.
VIM-VQ 5.004 4.256 6.915 0.892 0.128 0.861 0.601 0.101 1.109 0.055
VIM (Ours) 5.252 4.511 6.981 0.758 0.064 0.901 0.568 0.059 0.691 0.019
We conducted ablation studies comparing the VQ-VAE-based model with our RQ-VAE-based approach, as shown in Table C.2. The RQ-VAE-based motion tokenizer outperformed the VQ-VAE model in motion reasoning tasks, achieving higher scores in coherence, alignment, and naturalness. This improvement is attributed to reduced information loss, allowing our model to capture finer motion details while also enhancing its motion-to-text retrieval precision.
For generation and editing tasks, the VQ-VAE model achieved slightly better text-to-motion retrieval accuracy but performed worse in FID and MPJPE across editing, reaction generation, and T2M tasks, indicating degraded motion quality and less precise motion details. In contrast, our approach reduced MPJPE by 0.055 for reaction generation, preserving joint dynamics and producing more realistic and natural motions. VQ-VAE’s limitations are especially problematic for modeling interactive motions, where precise relative positioning is crucial, making its information loss and reconstruction quality more evident.
Figure C.1: Ablation Studies on codebook size and depth. We measured Top-3 retrieval-precision accuracy and FiD on the reconstructed motion.
We analyzed the effect of varying codebook size (128, 256, 512, 1024) and code depth (2, 3, 4, 5) on both reconstructed motion quality (measured by FID) and retrieval precision at top-3, and observed that increasing the codebook size from 128 to 512 reduced FID while simultaneously improving retrieval precision, indicating richer and more accurate motion representations, whereas moving to 1024 yielded diminishing returns at higher computational cost. Likewise, increasing the code depth from 2 to 4 provided better reconstruction quality and retrieval performance by allowing the model to capture more complex motion patterns, but further increasing the depth (e.g., to 5) showed marginal or even negative gains. Consequently, we selected 512 codes and a depth of 4 as the best trade-off between quality, retrieval accuracy, and efficiency.
D Expansion to Multi-Human Motion ($\geq 3$) Generation
Figure D.2: Our method demonstrates its extendability by generating multi-person interactions ($\geq 3$ people) through iterative prompting. Despite being trained on two-person scenarios, our framework conditions new motions on prior interactions, enabling the synthesis of natural group dynamics from textual descriptions.
Our method demonstrates the capability to extend motion generation to multi-person interactions ($\geq 3$ individuals) through iterative prompting, despite being trained exclusively on two-person motion scenarios. Specifically, we first generate motion for a pair of individuals based on the given text description, and then the motion of additional participants is synthesized while conditioning on the pre-existing interactions. This approach ensures that the newly generated motions remain coherent and contextually appropriate within the evolving group dynamics. We argue that this extendability is a key advantage of our versatile framework, as it allows for the scalable generation of complex human interactions without requiring additional multi-person training data. An example of this process is illustrated in Figure D.2, where the model successfully generates a realistic group conversation scene based on a textual prompt.
E Illustration on Data Collection Pipeline
Figure E.3: Overview of synthetic data generation for multi-turn conversations with interactive motions. (a) Motion captions and instructions are generated using GPT-4o based on interactions between two characters, followed by (b) the corresponding motion being synthesized using the InterGEN.
We illustrate the data collection pipeline for generating synthetic multi-turn conversations paired with interactive motions. As shown in Figure E.3, GPT-4-based prompts are used to create captions and instructions, which are then converted into corresponding motions using InterGEN [22].
F Additional Results for Traditional Motion Related Tasks
F.1 Motion to Text
Table F.3: Motion-to-Text
@@ F.1 Motion to Text
Table F.3 presents a comparative analysis of various methods on the motion-to-text generation task, focusing on retrieval precision and language evaluation metrics such as BLEU, METEOR, and Rouge-L. Among the task-specific approaches, TM2T achieves moderate retrieval precision scores, with a Top1 precision of 0.413, and language metrics of 0.192 for BLEU and 0.386 for METEOR. In contrast, our proposed unified method, VIM, attains significantly higher retrieval precision scores, with a Top1 precision of 0.669, and surpasses TM2T* in all language metrics, achieving a BLEU score of 0.230 and a METEOR score of 0.441. This indicates that VIM not only narrows but effectively reverses the performance gap between task-specific and unified approaches in this task. The superior performance of VIM in both retrieval precision and language generation metrics demonstrates its effectiveness in generating accurate and descriptive textual captions from motion inputs. By achieving higher scores than the task-specific TM2T, VIM showcases the potential of unified approaches to not only close but surpass the performance gap traditionally observed between task-specific and unified models in motion-to-text generation tasks. This advancement underscores the ability of VIM to balance motion understanding and language generation, leading to more coherent and relevant textual outputs.
F.2 Text to Motion
Table F.4: Text-to-Motion